100 (x-3) = 450-50x
Answer by checkley71(8403)

100(X-3)=450-50X
100X-300=450-50X
100X+50X=450+300
150X=750
X=750/150
X=5
PROOF
100(5-3)=450-50*5
100*2=450-250
200=200